OpenStudy (anonymous):

sq. rt of (x- 5) = 3

OpenStudy (lgbasallote):

square both sides to get rid of the square root...remember you want x...so... x-5 = 9 now isolate x x = 9+ 5 x = 14

OpenStudy (anonymous):

\[\sqrt{x-5} = 3\] \[x-5=9\] \[x=9+5\]